# Long time asymptotics for the KPII equation

### v3

@@ 4.1.1. Representation formulas of the Cauchy integrals
${\color[rgb]{0,0,0}=}$ ${\color[rgb]{0,0,0}\iint\frac{e^{2\pi itS_{0}(a;\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))}(2\pi i\xi_{1}^{\prime})^{j}\widetilde{s}_{c}(\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))}{p_{\lambda^{\prime}}(\xi^{\prime}_{1},\xi_{2}^{\prime})}\widetilde{f}(\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))d\xi^{\prime}_{1}d\xi^{\prime}_{2}},$
@@
$\mathcal{F}(t;\lambda^{\prime};x_{2}^{\prime};\xi_{1}^{\prime\prime})={\color[rgb]{0,0,0}\operatorname{sgn}(x_{2}^{\prime}+3t\lambda_{R}^{\prime})}\theta(-(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$
@@
(4.7) $\widetilde{\mathcal{C}T1}={\color[rgb]{0,0,0}\iint}[\,\frac{e^{+2\pi it(\pi^{2}{\xi^{\prime}_{1}}^{3}-\frac{3}{4}\frac{{\xi^{\prime}_{2}}^{2}}{\xi^{\prime}_{1}})}}{p_{\lambda^{\prime}}(\xi^{\prime}_{1},\xi^{\prime}_{2})}\,]^{\vee_{\xi^{\prime}_{1},\xi^{\prime}_{2}}}(ta-x_{1}^{\prime},-x_{2}^{\prime})[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x_{2}^{\prime},x^{\prime}_{2})dx_{1}^{\prime}dx_{2}^{\prime},$
@@
$\mathcal{F}^{(n)}(t;\lambda^{\prime};x_{2,n}^{\prime};\xi_{n}^{\prime\prime})={\color[rgb]{0,0,0}\operatorname{sgn}(x_{2,n}^{\prime}+3t\lambda_{R}^{\prime})}\theta(-(x_{2,n}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{n}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{n}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$
@@
$=$ ${\color[rgb]{0,0,0}-\frac{1}{2\pi i}}\iint\left[\frac{e^{\frac{2-\beta_{n-1}}{2}2\pi iS_{0}(\zeta^{\prime})}\left[\mathfrak{C}\mathfrak{T}\right]^{(n-1)}(t_{1},t_{2},t;\overline{\zeta}^{\prime})}{p_{\lambda^{\prime}}({\xi^{\prime}_{1}},{\xi^{\prime}_{2}})}\right]^{\vee_{{\xi^{\prime}_{1}},{\xi^{\prime}_{2}}}}(ta-x_{1,n}^{\prime},-x_{2,n}^{\prime})$
@@ 4.2. Long time asymptotics of $u_{2,0}(x)$ when $\color[rgb]{0,0,0}a>+\delta>0$
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
@@
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
@@ 4.3. Long time asymptotics of $u_{2,0}(x)$ when $\color[rgb]{0,0,0}a<-\delta<0$
Suppose the assumption of Theorem 1 is valid for $u_{0}\in\mathfrak{M}^{3,q}$. Then there exists a constant $C_{0}\gg 1$ such that, defining
@@
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
@@
$=$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|{\lambda^{\prime}}_{R}|-t^{-5/9})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-3.5/9})$
$+$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-1.9/9})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)$
$+$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|).$
From $\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})\neq 0$}, we can show that the $L^{1}(d\xi^{\prime\prime}_{n})$-norms of the amplitude function $\mathcal{F}^{(n)}$ over the corresponding domains for the first and second terms on the right-hand side of (4.68) are $o(t^{-1})$. Along with (2.3), the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$, and (2.12), the analysis reduces to studying the contribution over the domain corresponding to the third term, which is bounded by:
@@
$\times$ $\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)$
$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
@@
$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
@@
$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\theta(t^{-3.5/9}-||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|)\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
$\equiv$ $I_{1}+I_{2}+I_{3}.$
@@
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
@@ 5.1.1. Representation formulas of the Cauchy integrals
$\times$ $\int d\xi_{1}^{\prime\prime}e^{2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}(t;\lambda^{\prime};x_{1}^{\prime},x_{2}^{\prime};\xi_{1}^{\prime\prime})\equiv e^{i\pi tS_{0}(a;\lambda^{\prime})}\mathfrak{C}\mathfrak{T}_{1,(1)}1,$
@@
(5.4) $\left[\mathfrak{C}\mathfrak{T}1\right]^{1,(n)}(t_{1},t_{2},t;\lambda^{\prime})$
$=$ $\sum_{h=1}^{n}\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}\mathfrak{C}\mathfrak{T}_{1,(h)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{h}^{\prime\prime}),$
@@
(5.5) $\left[\mathfrak{C}\mathfrak{T}1\right]^{1,(n)}(t_{1},t_{2},t;\lambda^{\prime})$
$=$ $\sum_{h=1}^{n}\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}$
$\times$ $\{\,\mathfrak{C}\mathfrak{T}_{1,(h)}[1-\psi_{1,\xi_{h+1}^{\prime\prime}}(\xi_{h}^{\prime\prime})]+\mathfrak{C}\mathfrak{T}_{0,(h)}\psi_{1,\xi_{h+1}^{\prime\prime}}(\xi_{h}^{\prime\prime})\cdot{\color[rgb]{0,0,0}(2\pi i)}(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})\,\}$
$\times$ $\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{h}^{\prime\prime}),$
@@ 5.1.2. Asymptotics of the Cauchy integrals
Suppose the assumption of Theorem 1 is valid for $u_{0}\in\mathfrak{M}^{3,q}$, $\Xi(\lambda^{\prime})$ defined by (4.56), and $|a|>\delta>0$. Then, as $t\to\infty$,
@@ 5.2. Long time asymptotics of $u_{2,1}(x)$ when $\color[rgb]{0,0,0}a>+\delta>0$
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
@@
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
@@
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
@@
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
@@ 5.3. Long time asymptotics of $u_{2,1}(x)$ when $\color[rgb]{0,0,0}a<-\delta<0$
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
@@
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
@@
Suppose the assumption of Theorem 1 holds for $u_{0}\in\mathfrak{M}^{3,q}$. As $t\to\infty$,
@@
Assume the assumption of Theorem 1 holds for $u\in\mathfrak{M}^{3,q}$. As $t\to+\infty$,

### v4

@@ 4.1.1. Representation formulas of the Cauchy integrals
${\color[rgb]{0,0,0}=}$ ${\color[rgb]{0,0,0}-\iint\frac{e^{2\pi itS_{0}(a;\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))}(2\pi i\xi_{1}^{\prime})^{j}{\color[rgb]{0,0,0}\left[u_{0}m_{0}\right]^{\wedge}(\xi_{1},\xi_{2},\zeta)}}{p_{\lambda^{\prime}}(\xi^{\prime}_{1},\xi_{2}^{\prime})}\widetilde{f}(\zeta^{\prime}(\xi^{\prime}_{1},\xi^{\prime}_{2}))d\xi^{\prime}_{1}d\xi^{\prime}_{2}},$
@@
$\mathcal{F}(t;\lambda^{\prime};x_{2}^{\prime};\xi_{1}^{\prime\prime})={\color[rgb]{0,0,0}-\operatorname{sgn}(x_{2}^{\prime}+3t\lambda_{R}^{\prime})}\theta(-(x_{2}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{1}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{1}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$
@@
(4.7) $\widetilde{\mathcal{C}T1}={\color[rgb]{0,0,0}-\iint}[\,\frac{e^{+2\pi it(\pi^{2}{\xi^{\prime}_{1}}^{3}-\frac{3}{4}\frac{{\xi^{\prime}_{2}}^{2}}{\xi^{\prime}_{1}})}}{p_{\lambda^{\prime}}(\xi^{\prime}_{1},\xi^{\prime}_{2})}\,]^{\vee_{\xi^{\prime}_{1},\xi^{\prime}_{2}}}(ta-x_{1}^{\prime},-x_{2}^{\prime})[u_{0}\mathfrak{m}_{0}](x^{\prime}_{1}-\frac{2t_{2}}{3}x_{2}^{\prime},x^{\prime}_{2})dx_{1}^{\prime}dx_{2}^{\prime},$
@@
$\mathcal{F}^{(n)}(t;\lambda^{\prime};x_{2,n}^{\prime};\xi_{n}^{\prime\prime})={\color[rgb]{0,0,0}-\operatorname{sgn}(x_{2,n}^{\prime}+3t\lambda_{R}^{\prime})}\theta(-(x_{2,n}^{\prime}+3t\lambda^{\prime}_{R})(\xi_{n}^{\prime\prime}-\frac{\lambda^{\prime}_{I}}{2\pi})(\xi_{n}^{\prime\prime}+\frac{\lambda^{\prime}_{I}}{2\pi}))$
@@
$=$ ${\color[rgb]{0,0,0}-}\iint\left[\frac{e^{\frac{2-\beta_{n-1}}{2}2\pi iS_{0}(\zeta^{\prime})}\left[\mathfrak{C}\mathfrak{T}\right]^{(n-1)}(t_{1},t_{2},t;\overline{\zeta}^{\prime})}{p_{\lambda^{\prime}}({\xi^{\prime}_{1}},{\xi^{\prime}_{2}})}\right]^{\vee_{{\xi^{\prime}_{1}},{\xi^{\prime}_{2}}}}(ta-x_{1,n}^{\prime},-x_{2,n}^{\prime})$
@@ 4.2. Long time asymptotics of $u_{2,0}(x)$ when $\color[rgb]{0,0,0}a>+\delta>0$
Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
@@
Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
@@ 4.3. Long time asymptotics of $u_{2,0}(x)$ when $\color[rgb]{0,0,0}a<-\delta<0$
Suppose the assumption of Theorem 1 is valid. Then there exists a constant $C_{0}\gg 1$ such that, defining
@@
Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
@@
$=$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-1.9/9})\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})$
$+$ ${\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})}\theta(|\lambda_{R}^{\prime}|-t^{-5/9})\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9}).$
From $\color[rgb]{0,0,0}\psi_{r,r}(\lambda^{\prime}_{I})\neq 0$}, we can show that the $L^{1}(d\xi^{\prime\prime}_{n})$-norms of the amplitude function $\mathcal{F}^{(n)}$ over the corresponding domains for the first term on the right-hand side of (4.68) is $o(t^{-1})$. Along with (2.3), the factor $(\overline{\lambda}^{\prime}-\lambda^{\prime})$, and (2.12), the analysis reduces to studying the contribution over the domain corresponding to the second term, which is bounded by:
@@
$\times$ $\theta(t^{-1.9/9}-|\lambda_{R}^{\prime}|)\psi_{t^{-0.9/3},b}(\lambda_{I}^{\prime})\mathfrak{C}\mathfrak{T}_{0,(n)}\theta(t|\lambda_{R}^{\prime}|-|x_{2,n}^{\prime}|)\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
@@
$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|$
@@
$\times$ $\theta(||\xi_{n}^{\prime\prime}|-\frac{|\lambda^{\prime}_{I}|}{2\pi}|-t^{-7/9})\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(n-1)}|\equiv I_{1}+I_{2}+I_{3}.$
@@
Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
@@ 5.1.1. Representation formulas of the Cauchy integrals
(5.1) $\partial_{x_{1}}\widetilde{\mathcal{C}T1}(t_{1},t_{2},t;\lambda^{\prime})=e^{i\pi tS_{0}(a;\lambda^{\prime})}\iint dx_{1}^{\prime}dx_{2}^{\prime}\ \left(\partial_{x^{\prime}_{1}}[u_{0}\mathfrak{m}_{0}]\right)(x^{\prime}_{1}-\frac{2t_{2}}{3}x^{\prime}_{2},x^{\prime}_{2})e^{i\lambda^{\prime}_{I}(x_{1}^{\prime}+2\lambda^{\prime}_{R}x_{2}^{\prime})}$
$\quad\times\int d\xi_{1}^{\prime\prime}e^{2\pi it\mathfrak{G}^{\sharp}}\mathcal{F}(t;\lambda^{\prime};x_{1}^{\prime},x_{2}^{\prime};\xi_{1}^{\prime\prime})\equiv e^{i\pi tS_{0}(a;\lambda^{\prime})}\mathfrak{C}\mathfrak{T}_{1,(1)}1,$
@@
(5.4) $\left[\mathfrak{C}\mathfrak{T}1\right]^{1,(n)}(t_{1},t_{2},t;\lambda^{\prime})=\sum_{h=1}^{n}\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}\mathfrak{C}\mathfrak{T}_{1,(h)}\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{h}^{\prime\prime}),$
@@
(5.5) $\left[\mathfrak{C}\mathfrak{T}1\right]^{1,(n)}(t_{1},t_{2},t;\lambda^{\prime})=\sum_{h=1}^{n}\mathfrak{C}\mathfrak{T}_{0,(n)}\cdots\mathfrak{C}\mathfrak{T}_{0,(h+1)}$
$\quad\times\{\,\mathfrak{C}\mathfrak{T}_{1,(h)}[1-\psi_{1,\xi_{h+1}^{\prime\prime}}(\xi_{h}^{\prime\prime})]+\mathfrak{C}\mathfrak{T}_{0,(h)}\psi_{1,\xi_{h+1}^{\prime\prime}}(\xi_{h}^{\prime\prime})\cdot{\color[rgb]{0,0,0}(2\pi i)}(\xi_{h}^{\prime\prime}-\xi_{h+1}^{\prime\prime})\,\}$
$\quad\times\left[\mathfrak{C}\mathfrak{T}1\right]^{0,(h-1)}(t_{1},t_{2},t;{\lambda^{\prime}_{R}}+2\pi i\xi_{h}^{\prime\prime}),$
@@ 5.1.2. Asymptotics of the Cauchy integrals
Suppose the assumption of Theorem 1 is valid, $\Xi(\lambda^{\prime})$ defined by (4.56), and $|a|>\delta>0$. Then, as $t\to\infty$,
@@ 5.2. Long time asymptotics of $u_{2,1}(x)$ when $\color[rgb]{0,0,0}a>+\delta>0$
Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
@@
Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
@@
Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
@@
Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
@@ 5.3. Long time asymptotics of $u_{2,1}(x)$ when $\color[rgb]{0,0,0}a<-\delta<0$
Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
@@
Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
@@
Suppose the assumption of Theorem 1 holds. As $t\to\infty$,
@@
Assume the assumption of Theorem 1 holds. As $t\to+\infty$,